A 1200PF capacitor is charged by a 500v battery. It is disconnected
from the battery and is connected to 75MH inductor. Calculate 1) the initial
charge on the capacitor (ii) the maximum current (iii)the frequency and time
period of oscillation iv)total energy oscillating in the system. (3)
1
Expert's answer
2021-09-28T11:13:19-0400

The initial charge is


"Q=CV=0.6\space\mu\text{C}."

The maximum current:


"I_\text{max}=\frac{Q}{\sqrt{LC}}=0.063\text{ A}."

The period of oscillations:


"T=\frac 1f=2\pi\sqrt{LC}=5.96·10^{-5}\text{ s}."

The energy:


"E=\frac{Q^2}{2C}=375\text{ J}."